OpenStudy (christos):

Solve e^x - e^-x / 2 = 2

OpenStudy (anonymous):

Multiply both sides by e^x. It will be a quadratic equation.

OpenStudy (aravindg):

^^

OpenStudy (christos):

THANKS I WILL TRY!!

OpenStudy (christos):

What do I do after that?

OpenStudy (christos):

Help I am STUCK !!!!!!

OpenStudy (mertsj):

\[e^x-\frac{1}{2e^x}=2\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

I think \[e ^{x}-e ^{-x}\div2=2\] is the question

OpenStudy (mertsj):

Well then, that simplifies to: \[2e ^{2x}-4e ^{x}-1=0\]

OpenStudy (mertsj):

Which will not factor.

OpenStudy (christos):

I solved it boys, thanks a lot for your help! ! :)

OpenStudy (anonymous):

Let e^x be u and e^-x be 1/u Therefore, u-2/u=2 u^2-2=2u u^2-2u-2=0 We have, \[u=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] Now solve it.